# Computing Minimal Absent Words and Extended Bispecial Factors with CDAWG Space

### v1

@@ 6 Conclusions and Future Work
Our future work includes the following: Is it possible to extend our method for reporting length-bounded MAWs/EBFs of Section 4.3 to the problem of reporting MAWs/EBFs of length exactly $\ell$ for a query length $\ell>1$? This would be a space-improvement of the previous work [13]. Okabe et al. [21] proposed a (non-compacted) DAWG-based data structure that can report the symmetric difference of the sets of MAWs for multiple strings, using $O(n)$ space and time linear in the output size, where $n$ is the total length of the strings. Can our CDAWG-based data structure be extended to this problem?

### v2

@@ 6 Conclusions and Future Work
Our future work includes the following: Is it possible to extend our method for reporting length-bounded MAWs/EBFs of Section 4.3 to the problem of reporting MAWs/EBFs of length exactly $\ell$ for a query length $\ell>1$? This would be a space-improvement of the previous work [13]. Okabe et al. [21] proposed a (non-compacted) DAWG-based data structure that can report the symmetric difference of the sets of MAWs for multiple strings, using $O(n)$ space and time linear in the output size, where $n$ is the total length of the strings. Can our CDAWG-based data structures be extended to this problem?